40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 25 13:23:45 2021, Anchal, Summary, BHD, Before photos of BSC 
    Reply  Mon Oct 25 17:37:42 2021, Anchal, Summary, BHD, Part I of BHR upgrade - Removed optics from BSC rn_image_picker_lib_temp_44cb790a-c3b4-42aa-8907-2f9787a02acd.jpgrn_image_picker_lib_temp_0fd8f4fd-64ae-4ccd-8422-cfe929d4eeee.jpg
       Reply  Tue Oct 26 16:56:22 2021, Paco, Summary, BHD, Part I of BHR upgrade - Locked PMC and IMC 
          Reply  Wed Oct 27 16:27:16 2021, Anchal, Summary, BHD, Part II of BHR upgrade - Prep OMCNorthBefore.jpegOMCNorthAfter.jpeg
             Reply  Wed Oct 27 16:31:35 2021, Anchal, Summary, BHD, Part III of BHR upgrade - Removal of PR2 Small Suspension 
                Reply  Thu Oct 28 16:32:32 2021, Paco, Summary, BHD, Part IV of BHR upgrade - Removal of BSC eastern optics 
                   Reply  Fri Nov 5 12:21:16 2021, Anchal, Summary, BHD, Part VI of BHR upgrade - Removal of ITMYC optics XendFlowBenchLeftEnd.jpgXendFlowBenchMiddle.jpgXendFlowBenchRightEnd.jpg
                      Reply  Wed Dec 15 16:11:38 2021, Anchal, Summary, BHD, Part VIII of BHR upgrade - Placed LO1 signal-2021-12-15-161437.jpeg
                         Reply  Wed Dec 15 17:44:18 2021, Koji, Summary, BHD, Part VIII of BHR upgrade - Placed LO1 
                            Reply  Thu Jan 6 21:04:41 2022, Anchal, Summary, BHD, Part VIII of BHR upgrade - LO1 OSEMs inserted 
                               Reply  Mon Jan 10 14:52:51 2022, Anchal, Summary, BHD, LO1 OSEMs roughly calibrated and noise measured LO1_vs_ETMX_OSEM_Spectrum_LF_x30_Gain.pdfLO1_vs_ETMX_OSEM_Spectrum_HF_x30_Gain.pdf
                                  Reply  Mon Jan 10 18:36:41 2022, Anchal, Summary, BHD, LO1 free swinging test set to trigger 
                                     Reply  Tue Jan 11 12:19:12 2022, Anchal, Summary, BHD, LO1 Input Matrix Diagonalization performed. LO1_SUS_InpMat_Diagnolization.pdfLO1_FreeSwingData_PeakFitting.pdf
                         Reply  Thu Jan 6 11:54:20 2022, Anchal, Summary, BHD, Part IX of BHR upgrade - Placed AS1 and AS4 
                            Reply  Fri Jan 7 16:17:42 2022, Anchal, Summary, BHD, Part IX of BHR upgrade - Placed AS1 and AS4 filters 
                               Reply  Sat Jan 8 16:01:42 2022, Paco, Summary, BHD, Part IX of BHR upgrade - Placed LO2 filters 
                               Reply  Tue Jan 11 18:18:29 2022, Anchal, Summary, BHD, Attempted OSEM installation on AS1 
                                  Reply  Tue Jan 11 18:40:25 2022, Anchal, Summary, BHD, AS1 Sat Amp has a PCB issue signal-2022-01-11-184917.jpeg
                                     Reply  Thu Jan 13 09:48:41 2022, Anchal, Summary, BHD, AS1 Sat Amp fixed signal-2022-01-13-094823_001.jpegsignal-2022-01-13-094823_002.jpeg
                                  Reply  Thu Jan 13 12:24:08 2022, Anchal, Summary, BHD, AS1 SD and LR magnets broke 
                            Reply  Thu Jan 13 12:29:27 2022, Anchal, Summary, BHD, AS4 LR magnet broke 
                               Reply  Fri Jan 14 18:12:47 2022, Anchal, Summary, BHD, AS4 placed in ITMY Chamber, OSEMs connected 
                Reply  Thu Oct 28 17:01:54 2021, Anchal, Summary, BHD, Part III of BHR upgrade - Adding temp fixed flat mirror for PR2 
             Reply  Wed Oct 27 16:38:02 2021, Anchal, Summary, BHD, Part II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment 
                Reply  Wed Oct 27 18:11:37 2021, Koji, Summary, BHD, Part II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment 
                Reply  Wed Oct 27 18:16:45 2021, Koji, Summary, BHD, Part II of BHR upgrade - Relocation of TT2 and MMT1/2 alignment 
Message ID: 16425     Entry time: Mon Oct 25 17:37:42 2021     In reply to: 16424     Reply to this: 16429
Author: Anchal 
Type: Summary 
Category: BHD 
Subject: Part I of BHR upgrade - Removed optics from BSC 

[Anchal, Paco, Ian]


Clean room etiquettes

  • Two people in coverall suits, head covers, masks and AccuTech ultra clean gloves.
  • One person in just booties to interact with outside "dirty" world.
  • Anything that comes in chamber, first cleaned outside with clean cloth and IPA. Then cleaned by the "clean" folks. We followed this for allen keys, camera and beam finder card.
  • Once the chamber cover has been removed, cover the annulus with donut. We forgot to do this :(

Optics removal and changes

We removed the following optics from the BSC table and stored them in X-end flowbench with fan on. See attachment 1 and 2.

  1. IPPOS SM2
  2. GRX SM2
  3. PRM OL1
  4. PRMOL4
  5. IPPOS SM3
  6. IPANG SM1
  7. PRM OL2
  8. Unidentified optic inbetween IPPOS45P and IPPOS SM3
  9. Beam block behing PR3
  10. Beam block behind GR PBS
  11. GR PBS
  12. GRPERI1L (Periscope)
  13. PRMOL3
  14. IPPOS45P
  15. Cylindrical counterweight on North-west end of table.
  16. Cheap rectangular mirror on South west end of table (probably used for some camera, but not in use anymore)
  17. IPANGSM2

We also changed the direction of clamp of MMT1 to move it away from the center of the able (where PRM will be placed)

We screwed in the earthquake stops on PRM and BS from front face and top.

We unscrewed the cable post for BS and PRM oplevs and loved it in between SR3 and BS and screwed it lightly.

We moved the PRM, turned it anti-clockwise 90 degrees and brought it in between TT2 and BS. Now there is a clear line of sight between TT2 and PR2 on ITMY table.


Some next steps:

  • We align the input beam to TT2 by opening the "Injection Chamber" (formerly known as OMC chamber). While doing so, we'll clear unwanted optics from this table as well.
  • We open ITMX chamber, clear some POP optics. If SOS are ready, we would replace PR2 with SOS and put it in a new position.
  • Then we'll replace PR3 with an SOS and align the beam to BS.

These are next few days of work. We need atleast one SOS ready by Thursday.


Photos after today's work: https://photos.app.goo.gl/EE7Mvhw5CjgZrQpG6

Attachment 1: rn_image_picker_lib_temp_44cb790a-c3b4-42aa-8907-2f9787a02acd.jpg  5.572 MB  Uploaded Mon Oct 25 19:15:30 2021  | Hide | Hide all
rn_image_picker_lib_temp_44cb790a-c3b4-42aa-8907-2f9787a02acd.jpg
Attachment 2: rn_image_picker_lib_temp_0fd8f4fd-64ae-4ccd-8422-cfe929d4eeee.jpg  5.521 MB  Uploaded Mon Oct 25 19:15:41 2021  | Hide | Hide all
rn_image_picker_lib_temp_0fd8f4fd-64ae-4ccd-8422-cfe929d4eeee.jpg
ELOG V3.1.3-